OpenStudy (anonymous):

cot θ sin θ = 1/2 for 0 = θ < 2π

OpenStudy (ddcamp):

\[\cot(\theta) = \frac{ \cos(\theta) }{ \sin(\theta) }\]

OpenStudy (ddcamp):

So the question is asking when the cos of theta is one half.